# Learning a Neural Solver for Parametric PDEs to Enhance Physics-Informed Methods

### v3

@@ 1 Introduction
Partial Differential Equations (PDEs) are ubiquitous as mathematical models of dynamical phenomena in science and engineering. Solving PDEs is of crucial interest to researchers and engineers, leading to a huge literature on this subject (Evans, 2010; Salsa, 2015). Traditional approaches to solving PDEs such as finite difference, finite element analysis, or spectral methods (Zienkiewicz et al., 2005; LeVeque, 2007) often come with stability and convergence guarantees but suffer from a high computational cost. Improving numerical PDE solvers through faster and more accurate algorithms remains an active research topic (Zienkiewicz et al., 2005).
@@
PDE solvers usually rely on discretization and/or linearization of the problem through various techniques to simplify the computations. Iterative methods such as Jacobi, Gauss-Seidel, Conjugate Gradient, and Krylov subspace methods can then be used to solve the resulting systems. Unfortunately, many PDEs have an ill-conditioned nature, and these iterative processes can demand extensive computational resources. Preconditioning techniques are often essential to mitigate this, though they require precise customization to the specific PDE problem, making the development of effective solvers a significant research endeavor in itself. Yet, the computational demands, time, and expertise required to develop these algorithms sometimes make them impractical or sub-optimal for specific classes of problems. Instead of relying on hand-designed algorithms, researchers have investigated, as an alternative, the use of machine learning to train iterative PDE solvers Hsieh et al. (2019); Li et al. (2023a); Rudikov et al. (2024); Kopaničáková et al. (2023). These approaches usually parallel the classical numerical methods by solving a linear system resulting from the discretization of a PDE, for example, using finite differences or finite elements. A preconditioner is learned from data by optimizing a residual loss computed w.r.t. a ground truth solution obtained with a PDE solver. This preconditioner is used on top of a baseline iterative solver and aims at accelerating its convergence. Examples of baseline solvers are the conjugate gradient (Li et al., 2023a; Rudikov et al., 2024) or the Jacobi method Hsieh et al. (2019).
@@
Another recent research direction investigates the use of neural networks for building surrogate models in order to accelerate the computations traditionally handled by numerical techniques. These methods fall into two main categories: supervised and unsupervised. The supervised methodology consists of first solving the PDE using numerical methods to generate input and target data and then regressing the solution using neural networks in the hope that this surrogate could solve new instances of the PDE. Many models, such as Neural Operators, lie within this class (Li et al., 2020; Raonić et al., 2023; Bartolucci et al., 2023) and focus on learning the solution operator directly through a single neural network pass. Unsupervised approaches, involve considering a neural network as a solution of the PDE. The neural network parameters are found by minimizing the PDE residual with gradient descent. Methods such as Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019), or DeepRitz (E & Yu, 2018) fall under this category. This family of methods is attractive as it does not rely on any form of data, but only on information from the PDE residual. However, they exhibit severe difficulties during training (Krishnapriyan et al., 2021; Ryck et al., 2023), often requiring many optimization steps and sophisticated training schemes (Krishnapriyan et al., 2021; Rathore et al., 2024). The ill-conditioned nature of PDE residual loss appears again in this context, making standard optimizers such as Adam inappropriate (see section E.6 and section E.7 for a visualization of the ill conditioning of this loss landscape). A detailed review of the existing literature is described in appendix A.
@@
In this work, we consider having access to the PDE as in unsupervised approaches and also to some data for training our neural solver. Our objective is to solve the optimization issues mentioned above by learning an iterative algorithm that solves the PDE from its residual, defined as in the PINNs framework (see fig. 1). This neural solver is trained from data, either simulations or observations. Different from the classical ML training problem which aims at learning the parameters of a statistical model from examples, the problem we handle is learning to learn, Andrychowicz et al. (2016), i.e. learning an iterative algorithm that will allow us to solve a learning problem. When vanilla PINNs handle a single PDE instance, requiring retraining for each new instance, we consider the more complex setting of solving parametric PDEs, the parameters may include boundary or initial conditions, forcing terms, and PDE coefficients. Each specific instance of the PDE, sampled from the PDE parameter distribution, will then be considered as a training example. The objective is then to learn a solver from a sample of the parametric PDE distribution in order to accelerate inference on new instances of the equation. With respect to unsupervised approaches, our model implements an improved optimization algorithm, tailored to the parametric PDE problem at hand instead of using a hand-defined method such as stochastic gradient descent (SGD) or Adam. As demonstrated in the experimental section, this approach proves highly effective for the ill-posed problem of optimizing the PINNs objective, enabling convergence in just a few steps. This is further illustrated through gradient trajectory visualizations in section E.6. In the proposed methodology, the neural solver will make use of the gradient information computed by a baseline gradient method to accelerate its convergence. In our instantiation, we will use SGD as our baseline algorithm, but the method could be easily extended to other baselines. Our model deviates from the traditional preconditioning methods by directly optimizing the non-linear PDE residual loss (Raissi et al., 2019) without going through the discretization steps. Our contribution includes :
@@
  • Extending the comparison to several parametric PDEs with varying parameters from $1$d static to $2$d+time problems. We perform a comparison with baselines demonstrating a significant acceleration of the convergence w.r.t. baselines.
@@ 2 Motivation
Let us first motivate our objective with a simple example. Physics-informed neural networks (PINNs) are a promising tool for optimizing neural networks in an unsupervised way to solve partial differential equations (PDEs). However, these methods are notoriously difficult to train (Krishnapriyan et al., 2021; Ryck et al., 2023). As an illustrative example of this challenge, let us solve the zero-boundary Poisson equation in $1$d on $\Omega=[-\pi,\pi]$. Note that this section is intentionally informal, we provide rigorous statements and proofs in Appendix B.
@@
[…]
@@
As a simple example, consider the parametrization given by considering a linear combination of Fourier features widely used (Tancik et al., 2020) [^1] Note that even though the ansatz is linear in $\Theta$, it is not linear in $x$., $u_{\Theta}(x)=\sum_{k=-K}^{K}\theta_{k}\phi_{k}(x)$, with $\phi_{0}(x)=\tfrac{1}{\sqrt{2\pi}}$, $\phi_{-k}(x)=\tfrac{1}{\sqrt{\pi}}\cos(kx)$ and $\phi_{k}(x)=\tfrac{1}{\sqrt{\pi}}\sin(kx)$ for $1\leq k\leq K$.
@@
We believe that this simple example clearly illustrates and highlights the fact that PINNs–even when considering a linear basis, and when the PDE is linear–suffer heavily from ill-conditioning: if $500$ steps are required in order to achieve a given error when $K=5$, roughly speaking, 312 500312500312\>500312 500 steps are required for only $K=25$. This result extends to more general linear systems of equations and linear ansatz, as explained in appendix B.
@@ 3.2 Methodology
Physics-informed neural networks consider an ansatz $u_{\Theta}$ parametrized by some finite-dimensional $\Theta$. The parameters $\Theta$ are iteratively updated by minimizing a criterion $\mathcal{L}_{\textnormal{PDE}}$ (e.g. the PDE residual), which assesses how well the ansatz $u_{\Theta}$ meets the conditions specified in equations 6 and 7. As introduced for example for PINNs (Raissi et al., 2019) or Deep Galerkin method (Sirignano & Spiliopoulos, 2018), we consider $\mathcal{L}_{\textnormal{PDE}}$ to be given by the strong formulation of the residual $\mathcal{L}_{\textnormal{Res}}$, plus a boundary discrepancy term [^3] Note that other formulations of the loss may also be considered in a straightforward manner. $\mathcal{L}_{\textnormal{BC}}$: $\mathcal{L}_{\textnormal{PDE}}=\mathcal{L}_{{}_{\textnormal{Res}}}+\lambda\mathcal{L}_{\textnormal{BC}},\quad,\lambda>0$.
@@
As illustrated in section 2, performing gradient descent, or alternatives such as Adam and L-BFGS on such a highly ill-conditioned loss $\mathcal{L}_{\textnormal{PDE}}$ leads to severe training difficulties (Krishnapriyan et al., 2021). The key idea in our work is to improve a baseline gradient descent algorithms with the neural solver. More precisely starting from a baseline gradient algorithm, SGD in our instantiation, instead of considering the classical update, we first transform the gradient using a neural network $\mathcal{F}_{\varrho}$ with parameters $\varrho$, depending on the values of the PDE parameters as well as on other inputs such as the residual gradient provided by SGD: $\nabla_{\Theta}\mathcal{L}_{\textnormal{PDE}}$. The objective is to transform, through the neural solver $\mathcal{F}_{\varrho}$, the ill-conditioned problem into a new, simpler problem that requires fewer steps in order to achieve a given error.
Once the neural solver $\mathcal{F}_{\varrho}$ has been learned, inference can be performed on any new PDE as follows (see inference algorithm 1). Starting from an initial ansatz parameter $\Theta_{0}$, it is iteratively updated by this solver. At iteration $l$, the steepest direction of the loss $\mathcal{L}_{\textnormal{PDE}}$ is first computed with autograd. Then, the gradient is transformed, in a PDE parameter dependant way, with $\mathcal{F}_{\varrho}$:
@@
This approach can be seen as learning the iterates in a PDE solver to achieve a low loss, similar to the residual minimization methods in PDEs (Elman et al., 2014). By design, this solver is intended to be applicable to different PDEs, as well as various sources, boundary conditions, and initial conditions. This flexibility allows for a broad range of applications, making it a versatile tool in solving complex PDEs with varying characteristics.
Designed as a parametric PDE solver, $\mathcal{F}_{\varrho}$  [^4] In the following we use ”solver” with $\mathcal{F}_{\varrho}$ as a short-hand to refer to our proposed method. is trained with input target data from different sets of PDE parameters, as outlined in section 3.1. Once trained, it will be used without retraining on new instances of the PDE, i.e. with new values of the PDE parameters. The underlying hypothesis is that even though the solutions may be different for different inputs and parameters, the solution methodology remains relatively consistent. This consistency is expected to enhance the algorithm’s ability to generalize across novel scenarios effectively.
@@ 3.3 Training of a physics-informed solver
Choice of Ansatz $u_{\Theta}$. A very common choice (Shen et al., 2011) is to consider a family of basis functions $\Psi(x)=\{\psi_{i}(x)\}_{i=1}^{N}$ and consider the ansatz to be given by its linear span ${u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)}$. In the following, we consider this linear reconstruction, although our formulation is generic in the sense that it can also accommodate nonlinear variants.  [^5] Although we have found this may further complicate training.
As indicated in eq. 10, the solver $\mathcal{F}_{\varrho}$ will be trained from samples of the PDE parameter distribution $(\gamma,f,g)$ and from the associated samples of the solution $u$. We first describe below the inference step aiming at iteratively updating the parameters $\Theta$ of the solution function $u_{\Theta}$ while $\mathcal{F}_{\varrho}$ is held fixed. We then describe how the ${\varrho}$ parameters of the solver are trained. Please refer to fig. 1 that illustrates the interaction between the two steps.
@@
             Estimate $\Theta_{L}$ from $\Theta_{0},(\gamma,f,g)$ using algorithm 1
@@
             Update $\varrho$ with gradient descent from the data loss in eq. 10
@@
Inference The inference step is performed at fixed values of the solver parameters $\mathcal{F}_{\varrho}$. It consists, for a given instance of the PDE with parameters $(\gamma_{i},f_{i},g_{i})$, in finding the best $\Theta$ with a few steps of the solver using eq. 9. It is illustrated in fig. 1 - grey box and formalized in algorithm 1: starting from initial parameters $\Theta_{0}$, we compute the Physical loss $\mathcal{L}_{\textnormal{PDE}}$ using the ansatz $u_{\Theta_{0}}$. The PDE derivatives in $\mathcal{L}_{\textnormal{PDE}}$, can be computed by hand or automatic differentiation depending on the application  [^6] In our experiments, we computed the derivative by hand when possible since it fastens computations.. Then, eq. 9 is used to update the parameters $\Theta$ for a given number of steps $L$. The final solution is reconstructed using the linear combination, ${u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)}$, introduced in section 3.2 with the computed coefficients $\Theta_{L}$. Note that inference does not make use of the sampled target solutions $(u_{i})_{i=1}^{M}$ computed on the grid points$(x_{j})_{j=1}^{m}$.These targets are used exclusively for training the neural solver.
Training the neural solver Training amounts to learning the parameters of the solver $\mathcal{F}_{\varrho}$ and is performed with a training set of PDE parameters and simulation data considered as ground truth $(\gamma_{i},f_{i},g_{i},u_{i})_{i=1}^{M}$, corresponding to PDE instances (i.e. with different parameters $\gamma$ and/or forcing terms $f$ and/or initial/boundary conditions $g$). See fig. 1 - white box and algorithm 2. The objective is to learn a solver $\mathcal{F}_{\varrho}$ able, at inference, to converge to a target solution in a small ($2$ to $5$ in our experiments) number of steps. For that, an optimizer (Adam in our experiments) is used to update the $\mathcal{F}_{\varrho}$ parameters. The training algorithm makes use of the data associated with the different PDE instances by sampling PDEs in batches and running algorithm 1 on several PDE instances. For each PDE instance, one starts from an initial parameter value $\Theta_{0}$ and then performs two optimization steps (see algorithm 2): (i) one consists in solving in the ansatz parameters $\Theta$ using the neural solver using algorithm 1, leading to $\Theta_{L}$; (ii) the second one is the optimization of the solver parameters $\varrho$. We train the outputs directly to match the associated ground truth $(u_{i}(x_{j}))_{j=1}^{m}$ using the data loss:
$\mathcal{L}_{\text{DATA}}=\mathbb{E}_{\gamma,f,g}\left[||u_{\Theta_{L}}-u_{\gamma,f,g}||\right].$ (10)
The expectation is computed on the distribution of the PDE parameters $(\gamma,f,g)$. The solution $u$ is entirely determined by these parameters as indicated by the notation $u_{\gamma,f,g}$. $||u_{\Theta_{L}}-u||$ denotes a distance between the target $(u(x_{j}))_{j=1}^{m}$ and the forecast $(u_{\Theta_{L}}(x_{j}))_{j=1}^{m}$ with $m$ the trajectory size  [^7] To simplify the notation, we used a fixed grid size $m$. However, this framework can be used with different grid sizes, as well as irregular grids. See Ablation in appendix E, table 13. In practice, one samples a set of PDE instances $(\gamma_{i},f_{i},g_{i})$ and for each instance a corresponding sample $u_{i}$.
@@
Analyzing the behavior of the inference algorithm is challenging due to the non linear nature of the solver. We however could get some intuition using simplifying assumptions. We build on the ideas introduced in section 2 for the simple case of the Poisson equation, for which an explicit analytical solution could be derived. We provide in appendix B, a proof for a more general case and give below in theorem 1 our main result. This shows that the number of steps induced by $\mathcal{F}$ for the proposed algorithm is significantly less than the number of steps required by the baseline PINNs algorithm. This results is obtained under two main assumptions: (i) $\mathcal{F}$ behaves like its linearization and (ii) the descent operator $\mathcal{F}$ used in our algorithm, allows us to reach the optimum of $\mathcal{L}_{\text{DATA}}$.
@@
(Convergence rate in the linear case). Given a linear ansatz $u_{\Theta}(x)=\sum_{i=1}^{N}\theta_{i}\phi_{i}(x)$, assume the conditioner $\mathcal{F}$ behaves like its linearization $P=\text{Jacobian}(\mathcal{F})$, meaning that $\mathcal{F}$ can be replaced by $P$ at any point. Let A be the matrix derived from the PDE loss as eq. 3 for the Poisson equation or eq. 15 in the more general case. Denote by $\kappa(A)$ the condition number of the matrix $A$. The number of steps $N^{\prime}(\varepsilon)$ required to achieve an error $\|\Theta_{l}-\Theta^{\ast}\|_{2}\leq\varepsilon$ satisfies:
@@
We sketch the main insights here and refer to appendix B for the proof and a detailed analysis.
@@
  • In practice, and as shown in the experiments (section 4), the convergence rate is significantly improved w.r.t. the reference baseline gradient algorithm.

### v5

@@ 1 Introduction
Partial Differential Equations (PDEs) are ubiquitous as mathematical models of dynamical phenomena in science and engineering. Solving PDEs is of crucial interest to researchers and engineers, leading to a huge literature on this subject (PDE_evans10; PDEact_Salsa). Traditional approaches to solving PDEs such as finite difference, finite element analysis, or spectral methods (zienkiewicz2005finite; leveque2007finite) often come with stability and convergence guarantees but suffer from a high computational cost. Improving numerical PDE solvers through faster and more accurate algorithms remains an active research topic (zienkiewicz2005finite).
@@
PDE solvers usually rely on discretization and/or linearization of the problem through various techniques to simplify the computations. Iterative methods such as Jacobi, Gauss-Seidel, Conjugate Gradient, and Krylov subspace methods can then be used to solve the resulting systems. Unfortunately, many PDEs have an ill-conditioned nature, and these iterative processes can demand extensive computational resources. Preconditioning techniques are often essential to mitigate this, though they require precise customization to the specific PDE problem, making the development of effective solvers a significant research endeavor in itself. Yet, the computational demands, time, and expertise required to develop these algorithms sometimes make them impractical or sub-optimal for specific classes of problems. Instead of relying on hand-designed algorithms, researchers have investigated, as an alternative, the use of machine learning to train iterative PDE solvers hsieh2019learningNPDESconv; li23e-preconcgpde; rudikov2024fgcno; kopanivcakova2023enhancing. These approaches usually parallel the classical numerical methods by solving a linear system resulting from the discretization of a PDE, for example, using finite differences or finite elements. A preconditioner is learned from data by optimizing a residual loss computed w.r.t. a ground truth solution obtained with a PDE solver. This preconditioner is used on top of a baseline iterative solver and aims at accelerating its convergence. Examples of baseline solvers are the conjugate gradient (li23e-preconcgpde; rudikov2024fgcno) or the Jacobi method hsieh2019learningNPDESconv.
@@
Another recent research direction investigates the use of neural networks for building surrogate models in order to accelerate the computations traditionally handled by numerical techniques. These methods fall into two main categories: supervised and unsupervised. The supervised methodology consists of first solving the PDE using numerical methods to generate input and target data and then regressing the solution using neural networks in the hope that this surrogate could solve new instances of the PDE. Many models, such as Neural Operators, lie within this class (li2020fno; cno; reno) and focus on learning the solution operator directly through a single neural network pass. Unsupervised approaches, involve considering a neural network as a solution of the PDE. The neural network parameters are found by minimizing the PDE residual with gradient descent. Methods such as Physics-Informed Neural Networks (PINNs) (PINNs_Raissi19), or DeepRitz (DR1) fall under this category. This family of methods is attractive as it does not rely on any form of data, but only on information from the PDE residual. However, they exhibit severe difficulties during training (krishnapriyan_characterizing; deryckManu2023PreconditioningPINNs), often requiring many optimization steps and sophisticated training schemes (krishnapriyan_characterizing; RathoreICML2024). The ill-conditioned nature of PDE residual loss appears again in this context, making standard optimizers such as Adam inappropriate (see section˜E.6 and section˜E.7 for a visualization of the ill conditioning of this loss landscape). A detailed review of the existing literature is described in appendix˜A.
@@
In this work, we consider having access to the PDE as in unsupervised approaches and also to some data for training our neural solver. Our objective is to solve the optimization issues mentioned above by learning an iterative algorithm that solves the PDE from its residual, defined as in the PINNs framework (see fig.˜1). This neural solver is trained from data, either simulations or observations. Different from the classical ML training problem which aims at learning the parameters of a statistical model from examples, the problem we handle is learning to learn, andrychowicz2016l2lbygdbygd, i.e. learning an iterative algorithm that will allow us to solve a learning problem. When vanilla PINNs handle a single PDE instance, requiring retraining for each new instance, we consider the more complex setting of solving parametric PDEs, the parameters may include boundary or initial conditions, forcing terms, and PDE coefficients. Each specific instance of the PDE, sampled from the PDE parameter distribution, will then be considered as a training example. The objective is then to learn a solver from a sample of the parametric PDE distribution in order to accelerate inference on new instances of the equation. With respect to unsupervised approaches, our model implements an improved optimization algorithm, tailored to the parametric PDE problem at hand instead of using a hand-defined method such as stochastic gradient descent (SGD) or Adam. As demonstrated in the experimental section, this approach proves highly effective for the ill-posed problem of optimizing the PINNs objective, enabling convergence in just a few steps. This is further illustrated through gradient trajectory visualizations in section˜E.6. In the proposed methodology, the neural solver will make use of the gradient information computed by a baseline gradient method to accelerate its convergence. In our instantiation, we will use SGD as our baseline algorithm, but the method could be easily extended to other baselines. Our model deviates from the traditional preconditioning methods by directly optimizing the non-linear PDE residual loss (PINNs_Raissi19) without going through the discretization steps. Our contribution includes :
@@
  • Extending the comparison to several parametric PDEs with varying parameters from 11d static to 22d+time problems. We perform a comparison with baselines demonstrating a significant acceleration of the convergence w.r.t. baselines.
@@ 2 Motivation
Let us first motivate our objective with a simple example. Physics-informed neural networks (PINNs) are a promising tool for optimizing neural networks in an unsupervised way to solve partial differential equations (PDEs). However, these methods are notoriously difficult to train (krishnapriyan_characterizing; deryckManu2023PreconditioningPINNs). As an illustrative example of this challenge, let us solve the zero-boundary Poisson equation in 11d on $\Omega=[-\pi,\pi]$. Note that this section is intentionally informal, we provide rigorous statements and proofs in Appendix B.
@@
$\mathcal{L}_{\textnormal{PDE}}=\mathcal{L}_{{}_{\textnormal{Res}}}+\lambda\mathcal{L}_{\textnormal{BC}},\quad\mathcal{L}_{{}_{\textnormal{Res}}}=\int_{\Omega}\left|u_{\Theta}^{\prime\prime}(x)-f(x)\right|^{2}dx,$
@@
As a simple example, consider the parametrization given by considering a linear combination of Fourier features widely used (FF) [^1] Note that even though the ansatz is linear in $\Theta$, it is not linear in $x$., $u_{\Theta}(x)=\sum_{k=-K}^{K}\theta_{k}\phi_{k}(x)$, with $\phi_{0}(x)=\tfrac{1}{\sqrt{2\pi}}$, $\phi_{-k}(x)=\tfrac{1}{\sqrt{\pi}}\cos(kx)$ and $\phi_{k}(x)=\tfrac{1}{\sqrt{\pi}}\sin(kx)$ for $1\leq k\leq K$.
@@
We believe that this simple example clearly illustrates and highlights the fact that PINNs–even when considering a linear basis, and when the PDE is linear–suffer heavily from ill-conditioning: if $500$ steps are required in order to achieve a given error when $K=5$, roughly speaking, 312 500312\>500 steps are required for only $K=25$. This result extends to more general linear systems of equations and linear ansatz, as explained in appendix˜B.
@@ 3.2 Methodology
Physics-informed neural networks consider an ansatz $u_{\Theta}$ parametrized by some finite-dimensional $\Theta$. The parameters $\Theta$ are iteratively updated by minimizing a criterion $\mathcal{L}_{\textnormal{PDE}}$ (e.g. the PDE residual), which assesses how well the ansatz $u_{\Theta}$ meets the conditions specified in equations 6 and 7. As introduced for example for PINNs (PINNs_Raissi19) or Deep Galerkin method (Sirignano2018), we consider $\mathcal{L}_{\textnormal{PDE}}$ to be given by the strong formulation of the residual $\mathcal{L}_{\textnormal{Res}}$, plus a boundary discrepancy term [^3] Note that other formulations of the loss may also be considered in a straightforward manner. $\mathcal{L}_{\textnormal{BC}}$: $\mathcal{L}_{\textnormal{PDE}}=\mathcal{L}_{{}_{\textnormal{Res}}}+\lambda\mathcal{L}_{\textnormal{BC}},\quad,\lambda>0$.
@@
As illustrated in section˜2, performing gradient descent, or alternatives such as Adam and L-BFGS on such a highly ill-conditioned loss $\mathcal{L}_{\textnormal{PDE}}$ leads to severe training difficulties (krishnapriyan_characterizing). The key idea in our work is to improve a baseline gradient descent algorithms with the neural solver. More precisely starting from a baseline gradient algorithm, SGD in our instantiation, instead of considering the classical update, we first transform the gradient using a neural network $\mathcal{F}_{\varrho}$ with parameters $\varrho$, depending on the values of the PDE parameters as well as on other inputs such as the residual gradient provided by SGD: $\nabla_{\Theta}\mathcal{L}_{\textnormal{PDE}}$. The objective is to transform, through the neural solver $\mathcal{F}_{\varrho}$, the ill-conditioned problem into a new, simpler problem that requires fewer steps in order to achieve a given error.
Once the neural solver $\mathcal{F}_{\varrho}$ has been learned, inference can be performed on any new PDE as follows (see inference algorithm˜1). Starting from an initial ansatz parameter $\Theta_{0}$, it is iteratively updated by this solver. At iteration $l$, the steepest direction of the loss $\mathcal{L}_{\textnormal{PDE}}$ is first computed with autograd. Then, the gradient is transformed, in a PDE parameter dependant way, with $\mathcal{F}_{\varrho}$:
@@
This approach can be seen as learning the iterates in a PDE solver to achieve a low loss, similar to the residual minimization methods in PDEs (elman2014finite). By design, this solver is intended to be applicable to different PDEs, as well as various sources, boundary conditions, and initial conditions. This flexibility allows for a broad range of applications, making it a versatile tool in solving complex PDEs with varying characteristics.
Designed as a parametric PDE solver, $\mathcal{F}_{\varrho}$  [^4] In the following we use ”solver” with $\mathcal{F}_{\varrho}$ as a short-hand to refer to our proposed method. is trained with input target data from different sets of PDE parameters, as outlined in section˜3.1. Once trained, it will be used without retraining on new instances of the PDE, i.e. with new values of the PDE parameters. The underlying hypothesis is that even though the solutions may be different for different inputs and parameters, the solution methodology remains relatively consistent. This consistency is expected to enhance the algorithm’s ability to generalize across novel scenarios effectively.
@@ 3.3 Training of a physics-informed solver
Choice of Ansatz $u_{\Theta}$. A very common choice (SpectralMethodsShen) is to consider a family of basis functions $\Psi(x)=\{\psi_{i}(x)\}_{i=1}^{N}$ and consider the ansatz to be given by its linear span $u_{\Theta}(x)=\sum_{i=0}^{N}\theta_{i}\psi_{i}(x)$}. In the following, we consider this linear reconstruction, although our formulation is generic in the sense that it can also accommodate nonlinear variants.  [^5] Although we have found this may further complicate training.
As indicated in eq.˜10, the solver $\mathcal{F}_{\varrho}$ will be trained from samples of the PDE parameter distribution $(\gamma,f,g)$ and from the associated samples of the solution $u$. We first describe below the inference step aiming at iteratively updating the parameters $\Theta$ of the solution function $u_{\Theta}$ while $\mathcal{F}_{\varrho}$ is held fixed. We then describe how the $\varrho$} parameters of the solver are trained. Please refer to fig.˜1 that illustrates the interaction between the two steps.
@@
      Estimate $\Theta_{L}$ from $\Theta_{0},(\gamma,f,g)$ using algorithm˜1
@@
      Update $\varrho$ with gradient descent from the data loss in eq.˜10
@@
[…]
Training the neural solver Training amounts to learning the parameters of the solver $\mathcal{F}_{\varrho}$ and is performed with a training set of PDE parameters and simulation data considered as ground truth $(\gamma_{i},f_{i},g_{i},u_{i})_{i=1}^{M}$, corresponding to PDE instances (i.e. with different parameters $\gamma$ and/or forcing terms $f$ and/or initial/boundary conditions $g$). See fig.˜1 - white box and algorithm˜2. The objective is to learn a solver $\mathcal{F}_{\varrho}$ able, at inference, to converge to a target solution in a small ($2$ to $5$ in our experiments) number of steps. For that, an optimizer (Adam in our experiments) is used to update the $\mathcal{F}_{\varrho}$ parameters. The training algorithm makes use of the data associated with the different PDE instances by sampling PDEs in batches and running algorithm˜1 on several PDE instances. For each PDE instance, one starts from an initial parameter value $\Theta_{0}$ and then performs two optimization steps (see algorithm˜2): (i) one consists in solving in the ansatz parameters $\Theta$ using the neural solver using algorithm˜1, leading to $\Theta_{L}$; (ii) the second one is the optimization of the solver parameters $\varrho$. We train the outputs directly to match the associated ground truth $(u_{i}(x_{j}))_{j=1}^{m}$ using the data loss:
[…]
The expectation is computed on the distribution of the PDE parameters $(\gamma,f,g)$. The solution $u$ is entirely determined by these parameters as indicated by the notation $u_{\gamma,f,g}$. $u_{\Theta_{L}}-u||$ denotes a distance between the target $(u(x_{j}))_{j=1}^{m}$ and the forecast $(u_{\Theta_{L}}(x_{j}))_{j=1}^{m}$ with $m$ the trajectory size  [^7] To simplify the notation, we used a fixed grid size $m$. However, this framework can be used with different grid sizes, as well as irregular grids. See Ablation in appendix E, table 13. In practice, one samples a set of PDE instances $(\gamma_{i},f_{i},g_{i})$ and for each instance a corresponding sample $u_{i}$.
@@
Analyzing the behavior of the inference algorithm is challenging due to the non linear nature of the solver. We however could get some intuition using simplifying assumptions. We build on the ideas introduced in section˜2 for the simple case of the Poisson equation, for which an explicit analytical solution could be derived. We provide in appendix˜B, a proof for a more general case and give below in theorem˜1 our main result. This shows that the number of steps induced by $\mathcal{F}$ for the proposed algorithm is significantly less than the number of steps required by the baseline PINNs algorithm. This results is obtained under two main assumptions: (i) $\mathcal{F}$ behaves like its linearization and (ii) the descent operator $\mathcal{F}$ used in our algorithm, allows us to reach the optimum of $\mathcal{L}_{\text{DATA}}$.
@@
(Convergence rate in the linear case). Given a linear ansatz $u_{\Theta}(x)=\sum_{i=1}^{N}\theta_{i}\phi_{i}(x)$, assume the conditioner $\mathcal{F}$ behaves like its linearization $P=\text{Jacobian}(\mathcal{F})$, meaning that $\mathcal{F}$ can be replaced by $P$ at any point. Let A be the matrix derived from the PDE loss as eq.˜3 for the Poisson equation or eq.˜15 in the more general case. Denote by $\kappa(A)$ the condition number of the matrix $A$. The number of steps $N^{\prime}(\varepsilon)$ required to achieve an error $\|\Theta_{l}-\Theta^{\ast}\|_{2}\leq\varepsilon$ satisfies:
@@
We sketch the main insights here and refer to appendix˜B for the proof and a detailed analysis.
@@
  • In practice, and as shown in the experiments (section˜4), the convergence rate is significantly improved w.r.t. the reference baseline gradient algorithm.